OpenStudy (anonymous):

Find the GCF of the terms of the polynomial. 16x 3 + 28x 5 y

OpenStudy (unklerhaukus):

What do you get when you factorise the first term? \[16x^3=(\dots)\times(\dots)\times \dots\times(\dots)\times(\dots)\]

OpenStudy (unklerhaukus):

factorise it like this \(12x^2=2\times2\times3\times x\times x\)

OpenStudy (unklerhaukus):

@loveis130, can you factor the first term?

OpenStudy (anonymous):

If it helps you to understand factoring, think of the factor tree where you break down a number like 10 to 2 * 5